# Dynamic Slack Bus

### v1

@@ I Introduction
The objective of this letter is to show that the conventional slack bus model for power flow analysis is a special case of a more general dynamic formulation.
@@ II Slack Bus as a Set of Differential Equations
where we have assumed that $p_{h,s}$ depends only on the states $\boldsymbol{\sigma}$ and $p_{h,t}$ on both the state and their first time derivative. The latter dependency is what makes $p_{h,t}$ vanish in steady-state. Equation (6) indicates that a device has slack bus capability only if it has a power source ($p_{s,h}$), as expected. On the other hand, the transient power component, which is due to some form of stored energy ($p_{t,h}$), vanishes in steady state. The examples further elaborate on this point and discuss the role of the transient component in the relevant case of the synchronous machine.
@@ III-A Synchronous Machines
where $\tau_{e,h}^{\max}$ is the maximum electrical torque; $\omega_{n}$ is the nominal synchronous speed with, typically, $\omega_{n}=1$ pu; $\delta_{h}$ and $\omega_{h}$ are the angle and the rotor speed, respectively; $\omega_{s}$ is the angular speed of a common reference angular speed (typically, the angular speed of the center of inertia); $\Omega_{b}$ is the angular frequency base in rad/s; and parameters have usual meanings [2]. Hence, in this model, $\boldsymbol{\sigma}=[\delta_{h},\omega_{h}]^{T}$. The power injection of the machine into the grid is in the form of (8), where the two components of the power are:
@@ III-B Primary Frequency Control
where $R_{h}$ is the droop coefficient and $p^{o}_{m,h}$ is the turbine power set point. Thus, (12) can be rewritten as:
@@ III-C Automatic Generation Control
Its output signal is shared with the turbine governors of the synchronous machines and adjusts their power set point. Equation (16) is thus rewritten as:
@@ III-E Grid-Following Converters
where $r_{\rm f}$, $\ell_{\rm f}$, $c_{\rm f}$ are the resistance, inductance and capacitance (expressed in pu) of the ac filter of the converter; $g_{\rm dc}$ is an equivalent conductance that accounts for the losses of the converter; $c_{\rm dc}$ is the condenser utilized to filter the ripple on the converter dc voltage; $v_{h}$ and $i_{h}$ are the ac voltage and current, respectively; $p_{\rm dc}=v_{\rm dc}i_{\rm dc}$ is the dc power injected into the converter. From (28), one obtains that the GFL can provide slack bus capability in steady-state only if the dc side includes a controlled energy source. For example, if the dc current of the energy source is regulated through a droop frequency controller, one has:
@@
$p_{t,h}$ $=T_{\rm dc}i^{\prime}_{\rm dc}i_{\rm dc}+c_{\rm dc}\,v^{\prime}_{\rm dc}v_{\rm dc}+\ell_{\rm f}\,i_{h}^{\prime}i_{h}+c_{\rm f}\,v^{\prime}_{h}v_{h}\,.$
In steady state, the first equation of (30) resembles the distributed slack model (2). Moreover, similarly to the case of the passive load, the transient power $p_{t,h}$ is due to the energy stored in the dc and ac filters, and it decays in the order of milliseconds in case of a disturbance.
@@ III-F Grid-Forming Converters
from where one can deduce that the source and transient powers of the GFM (omitting for simplicity the terms due to dc and ac filters and losses which are similar to the ones shown for the GFL) are:
@@
IV Conclusions
The letter shows that the well-known distributed slack bus model can be interpreted as the steady-state version of a set of differential-algebraic equations. The dynamic version includes a transient component which has a role in the weak version of the proposed definition. A variety of examples applies the proposed framework to a variety of devices, including passive loads and electronic power converters with both grid-following and grid-forming controls. Future work will elaborate on the proposed framework and study whether it is possible to set up devices with dynamic slack bus capability where the slack variables do not converge, in steady-state, to the system frequency.

### v2

@@ I Introduction
The objective of this letter is to propose a general framework to define and classify all possible static and dynamic slack bus models. The letter also shows that the conventional slack bus model for power flow analysis is a special case of a more general dynamic formulation. The motivation for the proposed general framework is twofold: (i) provide a precise mathematical definition of slack bus; and (ii) facilitate a better understanding of the functioning of power system devices, in particular, those providing inertial response and frequency control.
@@ II Slack Bus as a Set of Differential Equations
where we have assumed that $p_{h,s}$ depends only on the states $\boldsymbol{\sigma}$ and $p_{h,t}$ on both the state and their first time derivative. The latter dependency is what makes $p_{h,t}$ vanish in steady-state. For example, matching (2) with (6), one has $p_{h,s}\equiv p^{o}_{h}+k_{g,h}\hat{\sigma}$ and $p_{h,t}\equiv 0$. Equation (6) indicates that a device has slack bus capability only if it has a power source ($p_{s,h}$), as expected. On the other hand, the transient power component, which is due to some form of stored energy ($p_{t,h}$), vanishes in steady state. The case study further elaborates on this point and discusses the role of the transient component in the relevant cases of synchronous machines, as well as of GFL and GFM converters.
@@ III-A Synchronous Machines
where $\tau_{e,h}^{\max}$ is the maximum electrical torque; $\tau_{m,h}$ is the mechanical torque; $\omega_{n}$ is the nominal synchronous speed with, typically, $\omega_{n}=1$ pu; $\delta_{h}$ and $\omega_{h}$ are the angle and the rotor speed, respectively; $\omega_{s}$ is the angular speed of a common reference angular speed (typically, the angular speed of the center of inertia); $\Omega_{b}$ is the angular frequency base in rad/s; and parameters have the usual meaning, namely, $M_{h}$ is the mechanical starting time; and $D_{h}$ is the damping [2]. Hence, in this model, $\boldsymbol{\sigma}=[\delta_{h},\omega_{h}]^{T}$. The power injection of the machine into the grid is in the form of (8), where the two components of the power are:
@@ III-B Primary Frequency Control
where $R_{h}$ is the droop coefficient, $T_{h}$ is the time constant of the primary frequency control, and $p^{o}_{m,h}$ is the turbine power set point. Thus, (12) can be rewritten as:
@@ III-C Automatic Generation Control
where $\xi$ and $K_{o}$ are the state variable and the gain, respectively, of the AGC integrator. The AGC output signal $\xi$ is shared with the turbine governors of the synchronous machines and adjusts their power set point. Equation (16) is thus rewritten as:
@@ III-E Grid-Following Converters
where $\zeta_{h}$, $k^{\rm p}_{h}$ and $k^{\rm i}_{h}$ are the integrator state, proportional gain, and integral gain, respectively, of the PLL loop filter; and $\Delta\hat{\omega}_{h}$ is the estimated frequency deviation with respect to the synchronous speed at the point of connection of the converter at the grid.
@@
where $r_{\rm f}$, $\ell_{\rm f}$, $c_{\rm f}$ are the resistance, inductance and capacitance (expressed in pu) of the ac filter of the converter; $g_{\rm dc}$ is an equivalent conductance that accounts for the losses of the converter; $c_{\rm dc}$ is the condenser utilized to filter the ripple on the converter dc voltage; $v_{h}$ and $i_{h}$ are the ac voltage and current, respectively; $v_{\rm dc}$ and $i_{\rm dc}$ are the dc voltage and current, respectively; and $p_{\rm dc}=v_{\rm dc}i_{\rm dc}$ is the dc power injected into the converter. From (28), one obtains that the GFL can provide slack bus capability in steady-state only if the dc side includes a controlled energy source. For example, if the dc current of the energy source is regulated through a droop frequency controller, one has:
@@
$p_{t,h}$ $=T_{\rm dc}i^{\prime}_{\rm dc}i_{\rm dc}+c_{\rm dc}\,v^{\prime}_{\rm dc}v_{\rm dc}+\ell_{\rm f}\,i_{h}^{\prime}i_{h}+c_{\rm f}\,v^{\prime}_{h}v_{h}\,,$
where $i^{o}_{\rm dc}$ is the initial steady-state value of the dc current; and $T_{\rm dc}$ and $R_{\rm dc}$ are the time constant and the droop coefficient of the primary frequency controller of the GFL. In steady state, the first equation of (30) resembles the distributed slack model (2). Moreover, similarly to the case of the passive load, the transient power $p_{t,h}$ is due to the energy stored in the dc and ac filters, and it decays in the order of milliseconds in case of a disturbance.
@@ III-F Grid-Forming Converters
where $p_{h}^{\rm ref}$ is the reference active power and $D_{\alpha}$ and $H_{\alpha}$ are the regulator time constant and integral deviation, respectively. from where one can deduce that the source and transient powers of the GFM (omitting for simplicity the terms due to dc and ac filters and losses which are similar to the ones shown for the GFL) are:
@@
IV Case Study
This section illustrates the proposed framework of dynamic slack bus through a numerical case study. With this aim, we consider the WSCC 9-bus test system and consider the following scenarios: (i) base case system with 3 synchronous machines, AVRs and turbine governors; (ii) and (iii) scenarios where the synchronous machines are substituted with three GFMs converters; and (iv) a scenario where the synchronous machines are substituted with three GFLs converters. For scenarios (ii) and (iii), two GFM models are considered, namely, REGFM_A1 [7] and REGFM_B1 [8], respectively. All scenarios include primary frequency and voltage control. For all scenarios, the contingency is a loss of 20% of the load consumption at bus 5 at $t=1$ s. Simulations are carried out with the software tool Dome [9].
Figure 1 shows the trajectory of the voltage phase angle at bus 1 for the four scenarios. The phase angles of the voltages at the other buses show a similar transient behavior. The initial operating condition is the same and is stable for all scenarios. The conventional synchronous machines, as well as the two models of GFMs have enough stored energy to compensate the power balance during the transient. On the other hand, in scenario (iv), the GFLs do not have enough stored energy to cope with the transient conditions and the simulation stops instants after $t=1$ s. This result was to be expected, as GFLs do not provide an inertial response.
Figure 1: Trajectories of the voltage phase angle at bus 1 ($\theta_{1}$) of the WSCC 9-bus system for different generation scenarios.
V Conclusions
The letter shows that the well-known distributed slack bus model can be interpreted as the steady-state version of a set of differential-algebraic equations. The dynamic version includes a transient component which has a role in the weak version of the proposed definition. A variety of examples applies the proposed framework to a variety of devices, including passive loads and electronic power converters with both grid-following and grid-forming controls. Future work will elaborate on the proposed framework and study whether it is possible to set up devices with dynamic slack bus capability where the slack variables do not converge, in steady-state, to the system frequency. The author also aims at utilizing the proposed mathematical framework to define new mechanisms for converter-based resources to provide power balance support to the grid.